10) How many molecules and atoms are in 3 moles of N,H?
Answer #1

1 molecule of N2H4 contains 2 nitrogen atoms and 4 hydrogen atom.

Total atoms = 6

1 mole N2H4 is 6.022*1023 molecules - (Avogadro number)

3 moles N2H4 = 3*6.022*1023 molecules = 18.066*1023 molecules

Number of atoms in 3 moles of N2H4 = 6*18.066*1023 = 108.396*1023 atoms= 1.08396*1025 atoms